A stone is dropped from the top of a cliff. The splash it makes when striking the water below is heard 4.1s later. The speed of sound in air is 343 m/s. How high is the cliff?
Answer #1

Let t1 be the time taken by the rock to touch the water and

t2 be the time taken by the sound to reach us.

Given that t1 + t2 = 4.1 s    -------------(1)

Velocity of sound v = 343 m/s

Let S be the distance tavelled by the rock.
          t2 = S/v = S/343 ........... (2)
from eq of motion, S = (1/2)g(t1)2    (initial speed u =0)
         t2 = [(1/2)(9.8 m/s2)(t1)2]/(343 m/s)
     = 0.014286(t1)2 ..........(3)
substitute the eq (3) in eq (1), we get
     0.014286(t1)2 + t1 - 4.1 = 0
On solving the above quadratic equation we get
                      t1 = 3.884 s

Hence distance covered by the rock is

                      S = (1/2)(9.8 m/s2)(3.884)2

                         = 73.92 m